# A multi-messenger hierarchical triple merger gravitational-wave event pair GW190514-GW190521 inside AGN J124942.3 + 344929

### v1

@@ I Introduction
GW190521 [1] is one of the most massive binary black hole merger observed during the first three LIGO-Virgo observing runs, with component masses of approximately $85\,M_{\odot}$ and $66\,M_{\odot}$, producing a remnant black hole of around $142\,M_{\odot}$, which can be classified as an intermediate mass black hole. GW190514 [2] is another binary black hole merger, with component masses of approximately $39\,M_{\odot}$ and $28\,M_{\odot}$, resulting in remnant black hole of around $67\,M_{\odot}$, which overlaps with the second mass range of GW190521. Moreover, there is significant spatial overlap between the sky localizations and distances of GW190514 and GW190521 (see Fig. 1), suggesting a plausible scenario where these two events are part of a hierarchical triple merger. In this scenario, GW190514 would represent the first merger, producing a remnant black hole that later paired with another black hole to form the hierarchical system responsible for GW190521 [3].
@@
In addition, an optical electromagnetic counterpart candidate, ZTF19abanrhr, was reported by Graham et al. [4], detected by the Zwicky Transient Facility approximately 34 days after the gravitational wave event GW190521. Evidence in favor of an association between the gravitational wave event and the electromagnetic counterpart was also reported in [5, 6] (although [7, 8]) This counterpart, confirmed as a flare within the active galactic nuclei (AGN) AGN J124942.3 + 344929, was at the 78% spatial contour for GW190521’s sky localization. Interestingly, GW190514’s 90% sky localization also coincides with the position of ZTF19abanrhr (see Fig. 1). This raises an intriguing question: Could both GW190514 and GW190521 be a hierarchical triple merger originating from AGN J124942.3 + 344929, with ZTF19abanrhr serving as their electromagnetic counterpart?
@@
Theoretical models suggest that AGN disks are promising environments for massive and hierarchical binary black hole mergers [9, 10, 11], consistent with the observed events GW190514 and GW190521. In such environments, electromagnetic emissions are expected, likely produced by shocks from accreting merger remnants interacting with the baryon-rich, high-density surroundings [12, 13, 14, 15]. Additionally, AGN disks could facilitate an excess of eccentric mergers, which may provide a possible explanation for the non-zero eccentricity and significant spin–orbit tilt of GW190521 if it originated in such an environment [16]. Moreover, the final mass of GW190514 is consistent with the best-matching eccentric template for GW190521’s mass distribution [3].
@@
Motivated by this, we propose a compelling scenario: GW190514, detected by the Advanced LIGO [17], represents a first generation black hole merger occurring within AGN J124942.3 + 344929, which left behind a remnant black hole that rapidly merged with another black hole, facilitated by interactions with the gas, leading to the formation of GW190521, detected by the Advanced LIGO and the Advanced Virgo [18]. Concurrently, the accreting remnant black hole from either GW190521 or GW190514 produced an optical flare ZTF19abanrhr, which was observed by the Zwicky Transient Facility [19] approximately 40 days after the merger [4, 20].
@@ II Methods
We explore the potential association between GW190514 and GW190521 as hierarchical triple merger in AGN J124942.3 + 344929, associated with ZTF19abanrhr, using data ($d_{1}$ and $d_{2}$ for the gravitational signals [^1] https://zenodo.org/record/5117702, and $s$ for the electromagnetic signal [^2] https://lasair-ztf.lsst.ac.uk/objects/ZTF19abanrhr/) from the GWTC-2.1 release [2] and ZTF [19]. This is a hypothesis-testing question. Our analysis employs a Bayesian criterion to assess whether signals observed from these separate data sets share a common origin. The association ($\mathcal{H}_{1}$) at least requires that i) the localizations of GW190514, GW190521, and ZTF19abanrhr must be consistent (see Fig. 1); and ii) the final mass of GW190514 should align with the second mass in GW190521. In contrary, the coincidence ($\mathcal{H}_{0}$) corresponds to the scenario where the three detections are associated by chance. As a result, the odds $\mathcal{O}^{1}_{0}$ between the hypotheses $\mathcal{H}_{1}$ and $\mathcal{H}_{0}$ is given by: $\mathcal{O}^{1}_{0}(d_{1},d_{2},s)=\mathcal{P}^{1}_{0}\,\mathcal{B}^{1}_{0}(d_{1},d_{2},s)$, where $\mathcal{P}^{1}_{0}$ is the prior odds; and $\mathcal{B}^{1}_{0}(d_{1},d_{2},s)=p(d_{1},d_{2},s|\mathcal{H}_{1})/p(d_{1},d_{2},s|\mathcal{H}_{0})$ is the Bayes factor defined by the ratio of the evidences of the hypotheses.
The derivations and computational framework presented as follows build upon the foundational methodology established in Morton et al. [5], who demonstrated the astrophysical association between GW190521 and flare ZTF19abanrhr as its electromagnetic counterpart. Here we investigate the association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr.
@@ II.1 Environmental effects
Stellar-mass binary black holes residing in AGN disks near the central supermassive black hole are influenced by various environmental effects. These effects can introduce biases in the parameter estimates of gravitational wave events, particularly altering the observed mass and distance of sources. Following the analysis in the previous literature [5], both effects impact the properties of the sources.
The first effect, related to the gravitational potential of the supermassive black hole, leads to gravitational redshift.
@@
where $R_{\rm s}$ represents the Schwarzschild radius of the supermassive black hole, and $r$ is the distance between the binary black hole and the supermassive black hole. The gravitational redshift is calculated assuming a non-spinning supermassive black hole, although supermassive black holes in AGN disks are typically expected to spin [22]. This assumption holds because the effect of spin becomes negligible at the distances relevant for binary black holes with respect to the supermassive black hole [23, 24, 25].
The second effect, which is related to the motion of the binary black hole as it orbits the supermassive black hole, results in a relativistic redshift,
@@
where $\gamma=(1-v^{2})^{-1/2}$ is the Lorentz factor, $v$ is the magnitude of the velocity, and $\phi$ is the viewing angle between the velocity and the line of sight in the observer frame. Assuming that the binary black hole is on a circular orbit around a non-spinning supermassive black hole [26], its velocity is
@@
The effective distance of the source is also influenced by various redshift effects, with the relativistic redshift component, induced by the aberration of gravitational waves, being squared in its impact on the observed distance [32].
@@
Consequently, the environmental effects of the supermassive black hole, such as gravitational and relativistic redshifts, can be accounted for by simply replacing the observed mass $m^{z}$ with the effective mass $m^{z,{\rm eff}}$, and the luminosity distance $D_{L}$ with the effective luminosity distance $D_{L}^{\rm eff}$ [5].
@@ II.2 Bayesian criterion
In order to assess whether it is more likely that the two gravitational wave events, GW190514 and GW190521 (as a hierarchical triple merger), as well as the electromagnetic transient ZTF19abanrhr, originate from a common source, or whether the three detections are associated by chance, we employ a Bayesian statistical analysis. Given three detections in the GW190514 data set $d_{1}$, the GW190521 data set $d_{2}$, and the ZTF19abanrhr data set $s$, the odds $\mathcal{O}^{1}_{0}$ between the association model $\mathcal{H}_{1}$ and the coincidence model $\mathcal{H}_{0}$ is given by:
@@
is the Bayes factor and $\mathcal{P}^{1}_{0}=p(\mathcal{H}_{1})/p(\mathcal{H}_{0})$ is the prior odds. The prior odds for the two models can be simplified as the product of two parts: 1) the odds for the electromagnetic transient associated with one of the two gravitational wave events is $\sim$1/131131/131 / 13 [8], obtained from the fact that there are approximately $\approx$13131313 events similar to ZTF19abanrhr in the Zwicky Transient Facility alert stream over the given observing epoch [4]; 2) the odds for any two gravitational wave events detected by LIGO and Virgo being a hierarchical triple merger is $\sim$1/1000110001/10001 / 1000 [33] corresponding to the proportion of hierarchical merger events within the total population of detectable mergers in the optimistic scenario [34]. As a result, we adopt a conservative estimate for the astrophysical prior odds $\mathcal{P}^{1}_{0}\sim 10^{-5}$.
@@
where $p(D_{L},\Omega,m^{z}|d_{i},\mathcal{H}_{i})$ is the posterior probability density for the gravitational wave data set $d_{i}$ and $p(D_{L},\Omega,m^{z}|\mathcal{H}_{i})=p(D_{L},\Omega|\mathcal{H}_{i})\,p(m^{z}|\mathcal{H}_{i})=p(D_{L}|\mathcal{H}_{i})\,p(\Omega|\mathcal{H}_{i})\,p(m^{z}|\mathcal{H}_{i})$ is the prior probability distribution used during the parameter estimation run by the LIGO-Virgo-KAGRA (LVK) Collaboration. The parameter estimation prior is uniform in the sky localization and mass, and in the square of the luminosity distance. The possible travel between the mergers is neglected [3]. The probability density $p(D_{L},\Omega|s,\mathcal{H}_{i})=p(D_{L}|s,\mathcal{H}_{i})\,p(\Omega|s,\mathcal{H}_{i})$ is obtained from the localization of its host AGN J124942.3 + 344929. The conversion in $p(D_{L}|s,\mathcal{H}_{i})$ between luminosity distance and redshift takes into account environmental effects, which are discussed in detail in the corresponding models below.
@@
where the prior $p(m^{z}|\mathcal{H}_{i})$ differs in the different models. In the association model $\mathcal{H}_{1}$, $p(m^{z}|\mathcal{H}_{1})=p^{\rm 1g}_{m_{\rm f}}(m^{z}|\mathcal{H}_{1})$ represents the final detector-frame mass distributions of first generation binary black hole mergers that are not a result of a previous merger in the AGN formation channel. In the coincidence model $\mathcal{H}_{0}$, $p(m^{z}|\mathcal{H}_{0})=p^{\rm 1g}_{m_{\rm f}}(m^{z}|\mathcal{H}_{0})$ and $p(m^{z}|\mathcal{H}_{0})=p^{\rm hier}_{m_{\rm 2}}(m^{z}|\mathcal{H}_{0})$ refer to the final mass distributions of first-generation binary black hole mergers and the second mass distributions of hierarchical binary black hole mergers, respectively, in the detector frame, for formation channels other than AGN disks.
@@ II.3 Association model
In the association model $\mathcal{H}_{1}$, the gravitational wave events GW190514 and GW190521, as a hierarchical triple merger, and the flare ZTF19abanrh share a common origin. The localization is fixed to the host AGN J124942.3 + 344929, and as a result, the observed mass and distance of the source are altered due to environmental effects. In this case, the gravitational wave posterior probability densities used in our work are the detector-frame effective primary mass $m^{z,{\rm eff}}$ and effective luminosity distance $D_{L}^{\rm eff}$. These two parameters are independently measured in gravitational wave detection; however, under the common origin hypothesis, they depend on the additional redshifts introduced by environmental effects.
The luminosity distance of the electromagnetic source is expressed as $p(D_{L}|s,\mathcal{H}_{1})\rightarrow p(D_{L}^{\rm eff}|s,\mathcal{H}_{1})$:
@@
the redshift $z_{\rm AGN}=0.438$ of AGN J124942.3 + 344929333https://skyserver.sdss.org/dr12/en/tools/explore/Summary.aspx?id=1237665128546631763. The expressions for $z_{\rm grav}(r)$ and $z_{\rm rel}(r,\phi)$ in equation (14) are the ones in equations (1) and (2), respectively. Here we adopt the prior $p(\phi)$ to be uniform in ${\rm cos}(\phi)$. The radial distance prior, $p(r)$, between the binary black hole and the supermassive black hole is predicted based on migration traps in AGN disks [35]. We assume that this radial distance is uniformly distributed between 24.5 and 331 Schwarzschild radii [35]. It is important to note that the precise location of migration traps can vary depending on the disk model, and in some cases, there may not be a migration trap at all [36, 37, 38]. However, our results are robust and are not significantly affected by these uncertainties [5].
@@ II.5 Astrophysical distributions
To obtain the astrophysical distributions, we utilize a phenomenological population model introduced in Li et al. [11], which allows for the rapid simulation of binary black hole mergers parametrized, including hierarchical mergers, within dynamical formation channels. For the first-generation binary black hole merger distributions, we adopt the results from the analysis of GWTC-3 by the LVK Collaboration [39]. In particular, the primary masses of binary black hole mergers that do not result from a previous merger follow a distribution described by a PowerLaw+Peak model. The parameters of the PowerLaw+Peak model used for the coincidence model $\mathcal{H}_{0}$ are consistent with the results of GWTC-3. For the association model $\mathcal{H}_{1}$, the mass power-law index aligns with predictions for a top-heavy population of merging black holes in AGN disks [40]. The dimensionless spin-magnitude distribution of first-generation black holes is uniformly assigned within a beta distribution. In the model $\mathcal{H}_{0}$, its parameters are obtained from the GWTC-3, while the parameters determining the distribution’s mean and variance are set to $1.5$ and $3.0$ in the model $\mathcal{H}_{1}$, reflecting the angular momentum exchange between binaries and the disk [41]. In both models, the spin tilt angles are randomly selected from an isotropic distribution across a sphere.

### v2

@@ I Introduction
GW190521 [1] is one of the most massive binary black hole (BBH) merger observed during the first three LIGO-Virgo observing runs, with component masses of approximately $85\,M_{\odot}$ and $66\,M_{\odot}$, producing a remnant black hole (BH) of around $142\,M_{\odot}$, which can be classified as an intermediate mass BH. GW190514 [2] is another BBH merger, with component masses of approximately $39\,M_{\odot}$ and $28\,M_{\odot}$, resulting in remnant BH of around $67\,M_{\odot}$, which overlaps with the second mass range of GW190521. Moreover, there is significant spatial overlap between the sky localizations and distances of GW190514 and GW190521 (see Fig. 1), suggesting a plausible scenario where these two events are part of a hierarchical triple merger. In this scenario, GW190514 would represent the first merger, producing a remnant BH that later paired with another BH to form the hierarchical system responsible for GW190521 [3].
@@
In addition, an optical electromagnetic (EM) counterpart candidate, ZTF19abanrhr, was reported by Graham et al. [4], detected by the Zwicky Transient Facility approximately 34 days after the gravitational wave (GW) event GW190521. Evidence in favor of an association between the GW event and the EM counterpart was also reported in [5, 6] (although [7, 8]) This counterpart, confirmed as a flare within the active galactic nuclei (AGN) AGN J124942.3 + 344929, was at the 78% spatial contour for GW190521’s sky localization. Interestingly, GW190514’s 90% sky localization also coincides with the position of ZTF19abanrhr (see Fig. 1). This raises an intriguing question: Could both GW190514 and GW190521 be a hierarchical triple merger originating from AGN J124942.3 + 344929, with ZTF19abanrhr serving as their EM counterpart?
@@
Theoretical models suggest that AGN disks are promising environments for massive and hierarchical BBH mergers [9, 10, 11, 12], consistent with the observed events GW190514 and GW190521. In such environments, EM emissions are expected, likely produced by shocks from accreting merger remnants interacting with the baryon-rich, high-density surroundings [13, 14, 15, 16]. Additionally, AGN disks could facilitate an excess of eccentric mergers, which may provide a possible explanation for the non-zero eccentricity and significant spin–orbit tilt of GW190521 if it originated in such an environment [17]. Moreover, the final mass of GW190514 is consistent with the best-matching eccentric template for GW190521’s mass distribution [3].
@@
Motivated by this, we propose a compelling scenario: GW190514, detected by the Advanced LIGO [18], represents a first generation BH merger occurring within AGN J124942.3 + 344929, which left behind a remnant BH that rapidly merged with another BH, facilitated by interactions with the gas, leading to the formation of GW190521, detected by the Advanced LIGO and the Advanced Virgo [19]. Concurrently, the accreting remnant BH from either GW190521 or GW190514 produced an optical flare ZTF19abanrhr, which was observed by the Zwicky Transient Facility [20] approximately 40 days after the merger [4, 21].
@@ II Methods
We explore the potential association between GW190514 and GW190521 as hierarchical triple merger in AGN J124942.3 + 344929, associated with ZTF19abanrhr, using data ($d_{1}$ and $d_{2}$ for the gravitational signals [^1] https://zenodo.org/record/5117702, and $s$ for the EM signal [^2] https://lasair-ztf.lsst.ac.uk/objects/ZTF19abanrhr/) from the GWTC-2.1 release [2] and ZTF [20]. This is a hypothesis-testing question. Our analysis employs a Bayesian criterion to assess whether signals observed from these separate data sets share a common origin. The association ($\mathcal{H}_{1}$) at least requires that i) the localizations of GW190514, GW190521, and ZTF19abanrhr must be consistent (see Fig. 1); and ii) the final mass of GW190514 should align with the second mass in GW190521. In contrary, the coincidence ($\mathcal{H}_{0}$) corresponds to the scenario where the three detections are associated by chance. As a result, the odds $\mathcal{O}^{1}_{0}$ between the hypotheses $\mathcal{H}_{1}$ and $\mathcal{H}_{0}$ is given by: $\mathcal{O}^{1}_{0}(d_{1},d_{2},s)=\mathcal{P}^{1}_{0}\,\mathcal{B}^{1}_{0}(d_{1},d_{2},s)$, where $\mathcal{P}^{1}_{0}$ is the prior odds; and $\mathcal{B}^{1}_{0}(d_{1},d_{2},s)=p(d_{1},d_{2},s|\mathcal{H}_{1})/p(d_{1},d_{2},s|\mathcal{H}_{0})$ is the Bayes factor defined by the ratio of the evidences of the hypotheses.
The derivations and computational framework presented as follows build upon the foundational methodology established in Morton et al. [5], who demonstrated the astrophysical association between GW190521 and flare ZTF19abanrhr as its EM counterpart. Here we investigate the association between GW190514 and GW190521 as a hierarchical triple merger associated with ZTF19abanrhr.
@@ II.1 Environmental effects
Stellar-mass BBHs residing in AGN disks near the central supermassive BH are influenced by various environmental effects. These effects can introduce biases in the parameter estimates of GW events, particularly altering the observed mass and distance of sources. Following the analysis in the previous literature [5], both effects impact the properties of the sources.
The first effect, related to the gravitational potential of the supermassive BH, leads to gravitational redshift.
@@
where $R_{\rm s}$ represents the Schwarzschild radius of the supermassive BH, and $r$ is the distance between the BBH and the supermassive BH. The gravitational redshift is calculated assuming a non-spinning supermassive BH, although supermassive BHs in AGN disks are typically expected to spin [23]. This assumption holds because the effect of spin becomes negligible at the distances relevant for BBHs with respect to the supermassive BH [24, 25, 26].
The second effect, which is related to the motion of the BBH as it orbits the supermassive BH, results in a relativistic redshift,
@@
where $\gamma=(1-v^{2})^{-1/2}$ is the Lorentz factor, $v$ is the magnitude of the velocity, and $\phi$ is the viewing angle between the velocity and the line of sight in the observer frame. Assuming that the BBH is on a circular orbit around a non-spinning supermassive BH [27], its velocity is
@@
The effective distance of the source is also influenced by various redshift effects, with the relativistic redshift component, induced by the aberration of GWs, being squared in its impact on the observed distance [33].
@@
Consequently, the environmental effects of the supermassive BH, such as gravitational and relativistic redshifts, can be accounted for by simply replacing the observed mass $m^{z}$ with the effective mass $m^{z,{\rm eff}}$, and the luminosity distance $D_{L}$ with the effective luminosity distance $D_{L}^{\rm eff}$ [5].
@@ II.2 Bayesian criterion
In order to assess whether it is more likely that the two GW events, GW190514 and GW190521 (as a hierarchical triple merger), as well as the EM transient ZTF19abanrhr, originate from a common source, or whether the three detections are associated by chance, we employ a Bayesian statistical analysis. Given three detections in the GW190514 data set $d_{1}$, the GW190521 data set $d_{2}$, and the ZTF19abanrhr data set $s$, the odds $\mathcal{O}^{1}_{0}$ between the association model $\mathcal{H}_{1}$ and the coincidence model $\mathcal{H}_{0}$ is given by:
@@
is the Bayes factor and $\mathcal{P}^{1}_{0}=p(\mathcal{H}_{1})/p(\mathcal{H}_{0})$ is the prior odds. The prior odds for the two models can be expressed as the product of two components: (1) the odds that the EM transient is associated with one of the two GW events, estimated to be $\sim 1/13$ [8], based on the fact that approximately 13 transients similar to ZTF19abanrhr were detected in the Zwicky Transient Facility alert stream over the relevant observing period [4]; (2) the odds that both mergers in the hierarchical triple system are found in the observed BBHs, estimated to be $\sim 1/1000$ [34, 35], corresponding to the expected fraction of hierarchical triple mergers among detectable events in optimistic scenarios. The resulting combined astrophysical prior odds is approximately $1/13{,}000$. For robustness, we adopt a slightly more conservative value of $10^{-5}$ in our analysis, noting that the $1/1000$ estimate represents an optimistic upper limit and the typical value is likely smaller. This choice (approximately eight times smaller) ensures that the preference for the proposed association remains significant even under stricter prior assumptions.
@@
where $p(D_{L},\Omega,m^{z}|d_{i},\mathcal{H}_{i})$ is the posterior probability density for the GW data set $d_{i}$ and $p(D_{L},\Omega,m^{z}|\mathcal{H}_{i})=p(D_{L},\Omega|\mathcal{H}_{i})\,p(m^{z}|\mathcal{H}_{i})=p(D_{L}|\mathcal{H}_{i})\,p(\Omega|\mathcal{H}_{i})\,p(m^{z}|\mathcal{H}_{i})$ is the prior probability distribution used during the parameter estimation run by the LIGO-Virgo-KAGRA (LVK) Collaboration. The parameter estimation prior is uniform in the sky localization and mass, and in the square of the luminosity distance. The possible travel between the mergers is neglected [3]. The probability density $p(D_{L},\Omega|s,\mathcal{H}_{i})=p(D_{L}|s,\mathcal{H}_{i})\,p(\Omega|s,\mathcal{H}_{i})$ is obtained from the localization of its host AGN J124942.3 + 344929. The conversion in $p(D_{L}|s,\mathcal{H}_{i})$ between luminosity distance and redshift takes into account environmental effects, which are discussed in detail in the corresponding models below.
@@
where the prior $p(m^{z}|\mathcal{H}_{i})$ differs in the different models. In the association model $\mathcal{H}_{1}$, $p(m^{z}|\mathcal{H}_{1})=p^{\rm 1g}_{m_{\rm f}}(m^{z}|\mathcal{H}_{1})$ represents the final detector-frame mass distributions of first generation BBH mergers that are not a result of a previous merger in the AGN formation channel. In the coincidence model $\mathcal{H}_{0}$, $p(m^{z}|\mathcal{H}_{0})=p^{\rm 1g}_{m_{\rm f}}(m^{z}|\mathcal{H}_{0})$ and $p(m^{z}|\mathcal{H}_{0})=p^{\rm hier}_{m_{\rm 2}}(m^{z}|\mathcal{H}_{0})$ refer to the final mass distributions of first-generation BBH mergers and the second mass distributions of hierarchical BBH mergers, respectively, in the detector frame, for formation channels other than AGN disks.
@@ II.3 Association model
In the association model $\mathcal{H}_{1}$, the GW events GW190514 and GW190521, as a hierarchical triple merger, and the flare ZTF19abanrh share a common origin. The localization is fixed to the host AGN J124942.3 + 344929, and as a result, the observed mass and distance of the source are altered due to environmental effects. In this case, the GW posterior probability densities used in our work are the detector-frame effective primary mass $m^{z,{\rm eff}}$ and effective luminosity distance $D_{L}^{\rm eff}$. These two parameters are independently measured in GW detection; however, under the common origin hypothesis, they depend on the additional redshifts introduced by environmental effects.
The luminosity distance of the EM source is expressed as $p(D_{L}|s,\mathcal{H}_{1})\rightarrow p(D_{L}^{\rm eff}|s,\mathcal{H}_{1})$:
@@
the redshift $z_{\rm AGN}=0.438$ of AGN J124942.3 + 344929333https://skyserver.sdss.org/dr12/en/tools/explore/Summary.aspx?id=1237665128546631763. The expressions for $z_{\rm grav}(r)$ and $z_{\rm rel}(r,\phi)$ in equation (14) are the ones in equations (1) and (2), respectively. Here we adopt the prior $p(\phi)$ to be uniform in ${\rm cos}(\phi)$. The radial distance prior, $p(r)$, between the BBH and the supermassive BH is predicted based on migration traps in AGN disks [36]. We assume that this radial distance is uniformly distributed between 24.5 and 331 Schwarzschild radii [36]. It is important to note that the precise location of migration traps can vary depending on the disk model, and in some cases, there may not be a migration trap at all [37, 38, 39]. However, our results are robust and are not significantly affected by these uncertainties [5].
@@ II.5 Astrophysical distributions
To obtain the astrophysical distributions, we utilize a phenomenological population model introduced in Li et al. [11], which allows for the rapid simulation of BBH mergers parametrized, including hierarchical mergers, within dynamical formation channels. For the first-generation BBH merger distributions, we adopt the results from the analysis of GWTC-3 by the LVK Collaboration [40]. In particular, the primary masses of BBH mergers that do not result from a previous merger follow a distribution described by a PowerLaw+Peak model. The parameters of the PowerLaw+Peak model used for the coincidence model $\mathcal{H}_{0}$ are consistent with the results of GWTC-3. For the association model $\mathcal{H}_{1}$, the mass power-law index aligns with predictions for a top-heavy population of merging BHs in AGN disks [41]. The dimensionless spin-magnitude distribution of first-generation BHs is uniformly assigned within a beta distribution. In the model $\mathcal{H}_{0}$, its parameters are obtained from the GWTC-3, while the parameters determining the distribution’s mean and variance are set to $1.5$ and $3.0$ in the model $\mathcal{H}_{1}$, reflecting the angular momentum exchange between binaries and the disk [42]. In both models, the spin tilt angles are randomly selected from an isotropic distribution across a sphere.